OpenStudy (anonymous):

Solve the following system of equations by elimination: 3x + 2y= 8 x + y= 2. A. x= -4 y= -2 B. x= 4 y= -2 C. x= 4 y= 1 D. x= 6 y= 3

OpenStudy (shamil98):

3x + 2y = 8 -2(x + y = 2) 3x + 2y = 8 -2x -2y = -2 ------------- x = 6 6 + y = 2 y = -4

OpenStudy (anonymous):

thats not one of the answers

OpenStudy (shamil98):

my bad forgot to multiply -2 by 2 x = 4 4 + y = 2 y = -2